Im confused on this question, Please help.

An aqueous solution is that 30.0 percent sulfuric acid (H2SO4) by mass has a density of 1.307g/mL. Determine the MOLALITY of the solution. Thank you in advance for your help. Very much appreciated!
Answer #1

Molality is defined as the number of moles of solute per Kg of the solvent.

Suppose there is 100 g of the aqueous solution.

Mass of H2SO4 in the solution = 30 g

Moles of H2SO4 = 30 / 98 = 0.306 moles

Mass of Water (Solvent) = 70 g

Therefore there is 0.306 moles of solute in 70 g of solvent

Moles of solute in 1000 g (1 Kg) of solvent = 0.306 / 0.07 = 4.373 moles

Hence Molality = 4.373 m